OpenStudy (anonymous):

Thanks for helping! I need to simplify this equation to a+bi form. (1/2(cos(72 degrees)+isin(72 degrees))^5 To solve I should do 1/2^5 and multiply 72 degrees by 5 right? Thanks for helping!

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @Nnesha @wio @abb0t @zepdrix @Whitemonsterbunny17 @mathmate @jagr2713 @iki @Mehek14 @ikram002p

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

multiply \(72\) by \(5\)

OpenStudy (anonymous):

so if I do that and simplify it is right?

OpenStudy (anonymous):

you also have to take \(\left(\frac{1}{2}\right)^5\)

OpenStudy (anonymous):

if by "simplify" you mean "evaluate the functions" then yes

OpenStudy (anonymous):

yup, that is what I meant

OpenStudy (anonymous):

@satellite73 Can you help me find the complex fifth roots of 5-5sqrt(3)*i?

OpenStudy (anonymous):

yeah first write \[5-5\sqrt{3}i\] in trig form do you now how to do that?

OpenStudy (anonymous):

r=sqrt((5)^2+(-5sqrt(3)^2)) for r

OpenStudy (anonymous):

and -5sqrt(3)/5=tan(theta) for theta

OpenStudy (anonymous):

and then I make a rcis(theta)?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yeah \(r=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

from there what do I do?

OpenStudy (anonymous):

did you find \(\theta\)?

OpenStudy (anonymous):

theta=-pi/3 Sorry for being a little late. My browser froze :(

OpenStudy (anonymous):

r=10

OpenStudy (anonymous):

yeah looks like you got it

OpenStudy (anonymous):

From there how do I fin the complex fifth roots?

OpenStudy (anonymous):

find*

OpenStudy (anonymous):

divide the angle by 5

OpenStudy (anonymous):

-pi/15

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do I get the 5th root of r?

OpenStudy (anonymous):

just say it \[\sqrt[5]{10}\]

OpenStudy (anonymous):

you can't really evaluate any of this, just write it in trig form

OpenStudy (anonymous):

Is there more to it?

OpenStudy (anonymous):

It asks me for the complex fifth roots

OpenStudy (anonymous):

there are five of them

OpenStudy (anonymous):

go around the circle again, then repeat

OpenStudy (anonymous):

so I add 2pik to the angle?

OpenStudy (anonymous):

right, the original angle then divide by 5 again

OpenStudy (anonymous):

or else you can divide the circle in to 5 equal parts, with \[-\frac{\pi}{15}\]as one of them

OpenStudy (anonymous):

each time I add +2pi, should I divide r by 5, or is that a one time thing?

OpenStudy (anonymous):

add \(2\pi\) to \(-\frac{\pi}{3}\) then divide that one by \(5\)

OpenStudy (anonymous):

then repeat

OpenStudy (anonymous):

oh, does r change at all?

OpenStudy (anonymous):

no it is going to be \(\sqrt[5]{10}\) each time

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh wait, i think your \(r\) is wrong, check it again

OpenStudy (anonymous):

So would the first one be \[\sqrt[5]{10}(cis(\pi/3))?\]

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):

Are you sure, I got 10 again

OpenStudy (anonymous):

\[\sqrt{5^2+(5\sqrt{3})^2}\]

OpenStudy (anonymous):

\[\sqrt{25+75}=10\]

OpenStudy (anonymous):

you are right sorry

OpenStudy (anonymous):

it's okay, you're the one helping me :) Would the first answer be \[\sqrt[5]{10}(cis(\pi/3))\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

it's fine he is still viewing the chat so just give him a minute or to

OpenStudy (anonymous):

oh sorry was away for a minute

OpenStudy (anonymous):

kool

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

ok you found \(r=10,\theta =-\frac{\pi}{3}\)

OpenStudy (anonymous):

Is the thing I wrote earlier correct?

OpenStudy (anonymous):

so \[5-5\sqrt{3}i=10\left(\cos(-\frac{\pi}{3})+i\sin(-\frac{\pi}{3})\right)\]

OpenStudy (anonymous):

you want the five fifth roots

OpenStudy (anonymous):

\[\sqrt[5]{10}(cis(\pi/3))?\]

OpenStudy (anonymous):

one is \[\sqrt[5]{10}\left(\cos(-\frac{\pi}{15})+i\sin(-\frac{\pi}{15})\right)\]

OpenStudy (anonymous):

then add \(2\pi\) to \(-\frac{\pi}{3}\) to get \(\frac{5\pi}{3}\)

OpenStudy (anonymous):

oh yeah , what you said

OpenStudy (anonymous):

divide by 5 and the next one is \[\sqrt[5]{10}\left(\cos(\frac{\pi}{3})+i\sin(\frac{\pi}{3})\right)\]

OpenStudy (anonymous):

without adding 2pi?

OpenStudy (anonymous):

not sure what you mean

OpenStudy (anonymous):

you divided without adding 2pik where k=1

OpenStudy (anonymous):

to theta

OpenStudy (anonymous):

we now have two fifth roots (3 more to go) one is \[\sqrt[5]{10}\left(\cos(-\frac{\pi}{15})+i\sin(-\frac{\pi}{15})\right)\] and the other is \[\sqrt[5]{10}\left(\cos(\frac{\pi}{3})+i\sin(\frac{\pi}{3})\right)\]

OpenStudy (anonymous):

So I can +4pi and divide by 5 for another one?

OpenStudy (anonymous):

oh no we did add \(2\pi\) to the original angle of \(-\frac{\pi}{3}\) to get \(\frac{5\pi}{3}\) then we divided that one by 5 to get \(\frac{\pi}{3}\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

or add \(2\pi\) to \(\frac{5\pi}{3}\) same thing

OpenStudy (anonymous):

when you add you get \(\frac{11\pi}{3}\) divide that by 5 and get \(\frac{11\pi}{15}\) for the third root

OpenStudy (anonymous):

oh kk

OpenStudy (anonymous):

and 2pi+11pi/15

OpenStudy (anonymous):

lather, rinse, repeat

OpenStudy (anonymous):

oh no careful!

OpenStudy (anonymous):

don't add \(2\pi\) to \(\frac{11\pi}{15}\) add it to \(\frac{11\pi}{3}\)

OpenStudy (anonymous):

oh right, my bad

OpenStudy (anonymous):

hold on a second, it is not a mystery why you are doing this both sine and cosine are periodic with period \(2\pi\) that means the angle is NOT unique

OpenStudy (anonymous):

so all the original numbers are the same \[5-5\sqrt{3}i=10(\cos(-\frac{\pi}{3})+i\sin(-\frac{\pi}{3}))=10(\cos(\frac{5\pi}{3})+i\sin(\frac{5\pi}{3}))=...\] then are all equal

OpenStudy (anonymous):

you are just expressing the original number in trig form with different values of \(\theta\) then you divide each of them by 5 to get the different fifth roots

OpenStudy (anonymous):

yup, so I just continue doing what we were just doing?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the theta of the roots would be -pi/15, pi/3, 11pi/15, 17pi/15, and 23pi/15?

OpenStudy (anonymous):

@satellite73